OpenStudy (anonymous):

A taxi charges $2.05 per ride and $0.20 for each mile, which can be written as F = $2.05 + $0.20m. How many miles can you travel in the cab and have the fare be less than $10? A.) 43 B.) 15 C.) 39 D.) 25

OpenStudy (mathstudent55):

Set the formula 2.05 + .2m less than 10 and solve the inequality for m.

OpenStudy (agent0smith):

Fare is $10 or less, so use F = 10. 10 = $2.05 + $0.20m subtract 2.05 from both sides, then divide both sides by 0.2 to find m, miles.

jimthompson5910 (jim_thompson5910):

Fare < $10 F < 10 2.05 + 0.20m < 10 ... Plug in F = 2.05 + 0.20m 0.20m < 10-2.05 0.20m < 7.95 m < ???

OpenStudy (anonymous):

it would be 25?

jimthompson5910 (jim_thompson5910):

no

jimthompson5910 (jim_thompson5910):

how would you solve 0.20m < 7.95

jimthompson5910 (jim_thompson5910):

in other words, how would you isolate m?

OpenStudy (anonymous):

7.95/0.20=39?

jimthompson5910 (jim_thompson5910):

7.95/0.20 = 39.75 so you're close so m < 39.75 you round down to get m < 39 (since 39 is less than 39.75, it works) So if F < 10, then m < 39 where m is some whole number of miles

OpenStudy (anonymous):

so 43

jimthompson5910 (jim_thompson5910):

no, 39 not sure how you got 43

jimthompson5910 (jim_thompson5910):

anything over 39 miles will yield a fare over $10